OpenStudy (anonymous):

**Please Help, for an online Exam** what expression would you substitute into the first equation? 3x + y = 8 -x - 2y = -10

OpenStudy (anonymous):

Write it more clearly please

OpenStudy (anonymous):

For the following system, if you isolated x, in the second equation to use the Substitution Method what expression would you substitute into the first equation? 3x + y = 8 -x - 2y = -10

OpenStudy (anonymous):

A. -2y + 10 B. 2y + 10 C. 2y - 10 D. -2y - 10

OpenStudy (anonymous):

multiply the bottom equation by -1, mins 2y from both sides and then substitute the equation into the top one.

OpenStudy (anonymous):

x=10-2y Subing in: 3(10-2y)+y=8 => y= 22/5 if my maths is right so x= 6/5

OpenStudy (anonymous):

So, is the answer D?

OpenStudy (anonymous):

no the answer is a

OpenStudy (anonymous):

Okay, THANK YOU SM!

OpenStudy (anonymous):

your welcome :)